OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? ? inches Part B How many miles are represented by 1/4 inch on the map? ? miles

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

for part A, please you have to solve this proportion: 2:25=x:60 so, applying the fundamental property of the peopoetion, we have: 25*x=2*60 so, x=120/25=24/5=4 4/5 inches for part B, we have to solve this proportion:

OpenStudy (michele_laino):

2:25= (1/4):x applying the fundamental property of the proportion, I get: 2*x=25/4 so x=...? please continue

OpenStudy (anonymous):

Is x = 3.125?

OpenStudy (michele_laino):

x=25/8=3 1/8

OpenStudy (anonymous):

Thanks !!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

What about part b?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino where r u?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

What about Part B @Michele_Laino?